A 100 kg football player is traveling at 11 m/s when he is stopped in 0.1 seconds. What was the magnitude of the force that stopped the player ?
1
Expert's answer
2020-09-21T08:31:44-0400

According to Newton's second law:


"F=ma=m\frac{\Delta v}{\Delta t},\\\space\\
F=100\frac{(0-11)}{0.1}=-11000\text{ N}."


Negative sign means that this was a stopping (opposing) force.